Solve (4 answers are : -1/2,1/2,-isqrt(3),isqrt(3))
4x^4+11x^2=3
The answers are above I just do not get the steps. Thanks !

Answer by stanbon(75887)

4x^4+11x^2=3
-----------------
It's a quadratic in x^2.
4x^2^2 + 11x^2-3 = 0
Modify:
4x^2^2+12x^2-x^2-3 = 0
Factor:
4x^2(x^2+3)-(x^2+3) = 0
---
(x^2+3)(4x^2-1) = 0
x^2 = sqrt(3) or x^2 = 1/4
x = +-sqrt(3) or x = +-(1/2)
=============================
Cheers,
Stan H.
=============